Use the table below to answer the following question(s). Fiberia Accessories, a clothing retailer, is planning...

  1. Use the table below to answer the following question(s).

    Fiberia Accessories, a clothing retailer, is planning to introduce a new line of sweaters as part of the winter collection for $65 with an inventory of 1500. The main selling season is 60 days between November and December. The store then sells the remaining units in a clearance sale at 65 percent discount. Out of the 60 main retail days, Fiberia sells the sweaters at full retail price for only 45 days, while giving a discount of 25 percent for the remaining 15 days. The demand functions a, and b are given as 79.5 and 1.1 respectively.
    Marked Down Pricing Model for
    Fiberia Accessories's new sweater
    Data
    Retail Price $65
    Inventory 1500
    Selling Season (days) 60
    Days at Full Retail 45
    Intermediate Markdown 25 percent
    Clearance Markdown 65 percent
    Demand Function
    A 79.5
    B 1.1


    Calculate the revenue for the clearance sales period.

    $23,400

    $18,921.09

    $17,105.16

    $48,871.88
Answer #1

Total number of units sold during the full retail sales period.

Demand Function: Q = a - b.P = 79.5 - 1.1P

At full retail price period, P = $65

, Q (per day) = 79.5 - 1.1 x 65 = 8

The period is of 45 days, so, total sales = 8 x 45 = 360

Calculate the total revenue for the new line of sweaters

Revenue of the full price period = 360 x $65 = $23,400